Express in kg: 3470 g Express in kg: 3470 g Express in kg: 3470 g

Express in kg:
3470 g
Express in kg:
3470 g
Express in kg:
3470 g

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that,
= 1 kg = 1000 g
Then,
= 1 g = (1/1000) kg
= 3470 g
= (3470/1000) kg
= (3470/100)
= 3.470 kg

Thanks